# Deep Synoptic Array Science: Searching for Long Duration Radio Transients with the DSA-110

### v2

@@ II.3 Candidate Post-Processing
The second major post-processing stage is to apply a three-stage Convolutional Neural Network (CNN) classifier to the image (centered on the TOA) to identify and rule out Radio Frequency Interference (RFI). The CNN is trained on a set of simulated point sources, simulated near- and far-field RFI (examples of RFI are shown in Appendix B), and field data sets of RFI and pulsar B0329+54 from NSFRB commissioning. The model consists of successive convolutional blocks with batch normalization, nonlinear activations, and max-pooling, followed by a small fully connected head with dropout; the CNN parameter learning rates are optimized with an Adaptive Moment Estimation (Adam; adam2014method) optimizer using a binary cross-entropy loss. After applying the trained model to the 3D image cube (position, time, frequency), the network outputs a single logit for each candidate which, after a sigmoid activation, we interpret as the RFI probability $p_{\rm rfi}$ (0 indicates likely astrophysical; 1 indicates likely RFI). Simulated injection tests and pulsar observations were used to verify classifier performance. Positively identified candidates are assigned a candidate name and metadata (RA, DEC, TOA, DM, width, and S/N) are saved to a json file. A candidate plot is saved and pushed as a Slack alert for human inspection; this shows the time and frequency averaged image, the S/N as a function of trial DM and width for the final candidate pixel, the dynamic spectrum from the final candidate pixel, and the de-dispersed time series from the final candidate pixel. Finally, a trigger is sent to the correlator nodes to dump voltage data for a $26.85$ second interval around the burst to disk for offline analysis and localization; voltage triggering is still being implemented. Code for the NSFRB pipeline, including imaging, data transfer, searching, and post-processing, is publicly available at https://github.com/dsa110/dsa110-nsfrb.
@@ III.3 Pulsar Observations
Figure 6: Summary of Pulsar B0329+54 Monitoring Observations. (Top) The S/N of all pulses measured from predicted Times-of-Arrival (TOAs) are in blue. Initial candidates above the $8.5\sigma$ threshold are in grey, remaining candidates after PSF-based reduction are in black, and final candidates after classification are in red. (Bottom right) The S/N distribution of all single-pulses above $3\sigma$ are in blue, while ‘non-detections’ below $3\sigma$ are in grey. The median of the $3\sigma$ distribution is indicated in red and compared to the predicted S/N in black. (Bottom left) The folded pulse profile (S/N-weighted histogram of full timeseries) with 1036 phase bins is shown in red while a histogram of peak TOAs is shown in blue. Note that the folded profile TOAs are not phase connected between days, nor are they phase connected to past observations (e.g. deller2019microarcsecond). The profile from the European Pulsar Network (EPN) database is shown in black (seiradakis1995pulsar, https://psrweb.jb.man.ac.uk/epndb). The TOA histogram is normalized such that the sum equals the mean TOA S/N, while the EPN and folded profiles are scaled to match the TOA histogram peak. The sampling time 134 ms, which limits the pulse profile resolution, is shown in phase units as black dashed lines.
@@ IV.1 Survey Strategy
Figure 7: Coverage and Rate Limits for the DN-GPS pilot survey. (top) Each sub-plot shows one of 6 passes of the Galactic Plane (dashed black line). Center coordinates for each 3.35 s pointing are shown in red. The positions of known pulsars and magnetars are shown as diamonds; non-pulsar LPRTs detected to date are circles. A sample of White Dwarf-M Dwarf binaries from 2021MNRAS.506.5201R are shown as triangles. (bottom) 95% Poissonian burst rate upper limits are indicated by the colorbar for each $3^{\circ}\times 3^{\circ}$ grid region as discussed in Section IV. Sources of interest covered by the DN-GPS search region are labeled; of these, only B0329+54 is confidently detected.
@@ IV.2 Astrometric Calibration
Astrometric errors for candidates from the DN-GPS are estimated by comparing measured positions of sources from the Radio Fundamental Catalogue (RFC) to their milliarcsecond scale Very Long Baseline Interferometric (VLBI) localizations (petrov2025radio). 16 RFC calibrators are selected as those that have 1.4 GHz flux measurements from the National Radio Astronomy Observatory (NRAO) VLA Sky Survey [^7] The NVSS survey is used because the VLA D-configuration synthesized beamsize ($\sim 46^{\prime\prime}$) is comparable to the DSA-110 core’s ($\sim 31^{\prime\prime}$). Therefore, unresolved NVSS sources are most likely unresolved to the DSA-110, making flux measurements within a synthesized beam comparable. (NVSS; condon1998nrao) and fall within the DN-GPS field (see Figure 7). We form $175\times 175$ pixel images around each calibrator transit (between 1301-30 images, each averaged over 3.35 s) and use the casatools and Astropy.WCS modules to compute the RA and declination of each image pixel based on the current telescope pointing. Each image is fit with a 2D Gaussian around the main lobe of the RFC source to measure its position and align the images with each other, averaging over the full image set to maximize the signal-to-noise. This ‘alignment solution’ - the number of image pixels by which the source transits between each 3.35 s-fringestopped image - is also used to align images in the 0.67 s and 10.06 s search pipelines to correct small deviations from the expected transit due to the time and UV-grid resolution.
@@ V Discussion
While no candidates are detected, this is the first targeted search for LPRTs at 1.4 GHz, and the burst rate limit can be used to constrain models for LPRT emission. In this section, we contrast the detectability of two predominant models: the isolated neutron star model from cooper2024beyond and the White Dwarf-M Dwarf model from qu2025magnetic.
@@ V.1 Relevant Noise Estimates for Long and Short Periods
In the following analysis, we assume the LPRT sources are ‘always active’ for simplicity; some sources have appeared active for only months at a time (e.g. hurley2022radio) while others have been active for decades (e.g. hurley2023long). Therefore, we leave a more detailed treatment of activity windows to future work.
@@ V.2 Constraints on the White Dwarf-M Dwarf Binary and Neutron Star Models
where $v_{\rm p}\lesssim 10^{6}\,$cm yr-1 is the velocity of plastic flow, $B_{S}\approx 10^{15}$ G is the surface magnetic field strength, $A_{\rm fp}\approx 10\,$km is the size of the crust ‘footprint” undergoing plastic motion, and $\theta_{\rm fp}$ is the co-latitude of the footprint. By asserting that $L_{\rm diss,NS}>25\sigma_{\rm srch}d^{2}{\rm BW}$ for an LPRT at distance $d$ within this model, and using the characteristic $B_{S}\gtrsim(3.2\times 10^{19}\,G)\sqrt{P\dot{P}/2}$, we can solve for the maximum distance $d_{\rm max,NS}(P,\dot{P})$ to which the DN-GPS could detect the LPRT. Figure 10 (top) shows this result with $v_{\rm pl}=10^{4}$ cm yr-1; we see that for $\dot{P}>10^{-13}$ and $P<R^{-1}_{\rm 95}=4.3$ min the search is complete to only 1 kpc. To the edge of the galaxy, approximately 3030\,kpc, the search is only sensitive to $\dot{P}>10^{-10}$. Based on the $\dot{P}$ range of detected LPRTs with $P<R^{-1}_{\rm 95}=4.3$ min, this indicates that the search is not sensitive to magnetar-LPRTs beyond 11\,kpc. In addition, cooper2024beyond indicate that the plastic flow velocity can reasonably range from $1<v_{\rm pl}<10^{6}$ cm yr-1, which is proportional to $L_{\rm diss,NS}$ and can significantly alter the horizon distance. Therefore, the search is not complete for magnetar-LPRT objects; we conclude that a population of ultra-long period magnetars may still be present within the Galactic Plane despite our non-detection.
In the qu2025magnetic White Dwarf-M Dwarf binary model, the companion star orbits within the White Dwarf’s magnetosphere, creating a voltage drop between the two stars via the unipolar inductor effect (Faraday disk). As charge flows between the stars due to the potential drop they create an electrical current, and the magnetosphere dissipates power via Ohmic heating, driving pair creation. Pair charges then radiate via the relativistic electron cyclotron maser mechanism. We adjust the derivation from qu2025magnetic to allow non-zero spin-down $\dot{P}$; a brief derivation is provided in Appendix A, obtaining the following dissipation luminosity:
@@
Figure 10: $P-\dot{P}$ Constraints from the DN-GPS Survey Using a Neutron Star Model (top; cooper2024beyond) and a White Dwarf Model (bottom; qu2025magnetic). In each plot the thick red lines bound the allowed emission ranges for each model. Vertical grey lines indicate hard limits from the DN-GPS search due to the sampling time and Poissonian burst rate. Pulsars (diamonds), WD binary LPRTs (squares) and all other LPRTs (circles) are colored with their distance estimates, while those without distances are in red. The colorbar also highlights the maximum detectable distance ($d_{\rm max}(P,\dot{P})$) for each model (see Section IV) given the DN-GPS $25\sigma_{\rm srch,P<5\,min}\approx 1030$ mJy and $25\sigma_{\rm srch,P>5\,min}\approx 1200$ mJy detection limits for P<2P<2\,min and $P>2$ min, respectively. Red dashed lines are contours of constant $d_{\rm max}(P,\dot{P})$).
@@ VII Data Availability
The authors would like to thank Adam Dong, Maura McLaughlin, Duncan Lorimer, Scott Ransom, Bing Zhang, Tony Rodriguez, Assaf Horesh, Jayanta Roy, Bhaswati Bhattacharyya, Emmanuel Fonseca, Graham Doskotch, Natasha Hurley-Walker, and Kshitij Bane for creative discussions exploring the unique nature of LPRTs and novel ways to search for them. This work includes data from the Radio Fundamental Catalogue (RFC; www.doi.org/10.25966/dhrk-zh08) This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2139433. The authors thank staff members of the Owens Valley Radio Observatory and the Caltech radio group, whose tireless efforts were instrumental to the success of the DSA-110. The DSA-110 is supported by the National Science Foundation Mid-Scale Innovations Program in Astronomical Sciences (MSIP) under grant AST-1836018.
@@ Appendix A Derivation of Time-dependent Dissipation Luminosity for the White Dwarf Model
In this appendix, we derive equation 4 incorporating time dependence. Starting from Equation 9 in qu2025magnetic, the total voltage when the White Dwarf and M-dwarf are at a separation $a(t)$ is given by:
@@
Note that in general, we expect $\tau\ll\tau_{\rm age}\sim 1$ Gyr; for example, GPM J1839-10 was detected in archival data over the course of 36 years. Using its upper limit $\dot{P}<3.6\times 10^{-13}$ and measured P=31482.4P=31482.4\,s, we find $\dot{P}\tau/P\sim 10^{-8}$ for $\tau=36\,$years (horvath2025unified; men2025highly). Even for $\tau=1\,$Gyr, we have $\dot{P}\tau/P\sim 10^{-1}$. Therefore we conclude the solution is valid in general for the LPRT case. Note also that $\zeta$ can vary over a wide range ($\zeta\sim 22$ for GPM J1839-10), but we normalize it to unity for this discussion as in qu2025magnetic.

### v3

@@ II.3 Candidate Post-Processing
The second major post-processing stage is to apply a three-stage Convolutional Neural Network (CNN) classifier to the image (centered on the TOA) to identify and rule out Radio Frequency Interference (RFI). The CNN is trained on a set of simulated point sources, simulated near- and far-field RFI (examples of RFI are shown in Appendix B), and field data sets of RFI and pulsar B0329+54 from NSFRB commissioning. The model consists of successive convolutional blocks with batch normalization, nonlinear activations, and max-pooling, followed by a small fully connected head with dropout; the CNN parameter learning rates are optimized with an Adaptive Moment Estimation (Adam; Adam and others, 2014) optimizer using a binary cross-entropy loss. After applying the trained model to the 3D image cube (position, time, frequency), the network outputs a single logit for each candidate which, after a sigmoid activation, we interpret as the RFI probability $p_{\rm rfi}$ (0 indicates likely astrophysical; 1 indicates likely RFI). Simulated injection tests and pulsar observations were used to verify classifier performance. Positively identified candidates are assigned a candidate name and metadata (RA, DEC, TOA, DM, width, and S/N) are saved to a json file. A candidate plot is saved and pushed as a Slack alert for human inspection; this shows the time and frequency averaged image, the S/N as a function of trial DM and width for the final candidate pixel, the dynamic spectrum from the final candidate pixel, and the de-dispersed time series from the final candidate pixel. Finally, a trigger is sent to the correlator nodes to dump voltage data for a $26.85$ second interval around the burst to disk for offline analysis and localization; voltage triggering is still being implemented. Code for the NSFRB pipeline, including imaging, data transfer, searching, and post-processing, is publicly available at https://github.com/dsa110/dsa110-nsfrb.
@@ III.3 Pulsar Observations
Figure 6: Summary of Pulsar B0329+54 Monitoring Observations. (Top) The S/N of all pulses measured from predicted Times-of-Arrival (TOAs) are in blue. Initial candidates above the $8.5\sigma$ threshold are in grey, remaining candidates after PSF-based reduction are in black, and final candidates after classification are in red. (Bottom right) The S/N distribution of all single-pulses above $3\sigma$ are in blue, while ‘non-detections’ below $3\sigma$ are in grey. The median of the $3\sigma$ distribution is indicated in red and compared to the predicted S/N in black. (Bottom left) The folded pulse profile (S/N-weighted histogram of full timeseries) with 1036 phase bins is shown in red while a histogram of peak TOAs is shown in blue. Note that the folded profile TOAs are not phase connected between days, nor are they phase connected to past observations (e.g. Deller et al., 2019). The profile from the European Pulsar Network (EPN) database is shown in black (Seiradakis et al., 1995, https://psrweb.jb.man.ac.uk/epndb). The TOA histogram is normalized such that the sum equals the mean TOA S/N, while the EPN and folded profiles are scaled to match the TOA histogram peak. The sampling time 134 ms, which limits the pulse profile resolution, is shown in phase units as black dashed lines.
@@ IV.1 Survey Strategy
Figure 7: Coverage and Rate Limits for the DN-GPS pilot survey. (top) Each sub-plot shows one of 6 passes of the Galactic Plane (dashed black line). Center coordinates for each 3.35 s pointing are shown in red. The positions of known pulsars and magnetars are shown as diamonds; non-pulsar LPRTs detected to date are circles. A sample of White Dwarf-M Dwarf binaries from Rebassa-Mansergas et al. (2021) are shown as triangles. (bottom) 95% Poissonian burst rate upper limits are indicated by the colorbar for each $3^{\circ}\times 3^{\circ}$ grid region as discussed in Section IV. Sources of interest covered by the DN-GPS search region are labeled; of these, only B0329+54 is confidently detected.
@@ IV.2 Astrometric Calibration
Astrometric errors for candidates from the DN-GPS are estimated by comparing measured positions of sources from the Radio Fundamental Catalogue (RFC) to their milliarcsecond scale Very Long Baseline Interferometric (VLBI) localizations (Petrov and Kovalev, 2025). 16 RFC calibrators are selected as those that have 1.4 GHz flux measurements from the National Radio Astronomy Observatory (NRAO) VLA Sky Survey [^7] The NVSS survey is used because the VLA D-configuration synthesized beamsize ($\sim 46^{\prime\prime}$) is comparable to the DSA-110 core’s ($\sim 31^{\prime\prime}$). Therefore, unresolved NVSS sources are most likely unresolved to the DSA-110, making flux measurements within a synthesized beam comparable. (NVSS; Condon et al., 1998) and fall within the DN-GPS field (see Figure 7). We form $175\times 175$ pixel images around each calibrator transit (between 1301-30 images, each averaged over 3.35 s) and use the casatools and Astropy.WCS modules to compute the RA and declination of each image pixel based on the current telescope pointing. Each image is fit with a 2D Gaussian around the main lobe of the RFC source to measure its position and align the images with each other, averaging over the full image set to maximize the signal-to-noise. This ‘alignment solution’ - the number of image pixels by which the source transits between each 3.35 s-fringestopped image - is also used to align images in the 0.67 s and 10.06 s search pipelines to correct small deviations from the expected transit due to the time and UV-grid resolution.
@@ V Discussion
While no candidates are detected, this is the first targeted search for LPRTs at 1.4 GHz, and the burst rate limit can be used to constrain models for LPRT emission. In this section, we contrast the detectability of two predominant models: the isolated neutron star model from Cooper and Wadiasingh (2024) and the White Dwarf-M Dwarf model from Qu and Zhang (2025).
@@ V.1 Relevant Noise Estimates for Long and Short Periods
In the following analysis, we assume the LPRT sources are ‘always active’ for simplicity; some sources have appeared active for only months at a time (e.g. Hurley-Walker et al., 2022) while others have been active for decades (e.g. Hurley-Walker et al., 2023). Therefore, we leave a more detailed treatment of activity windows to future work.
@@ V.2 Constraints on the White Dwarf-M Dwarf Binary and Neutron Star Models
where $v_{\rm p}\lesssim 10^{6}\,$cm yr-1 is the velocity of plastic flow, $B_{S}\approx 10^{15}$ G is the surface magnetic field strength, $A_{\rm fp}\approx 10\,$km is the size of the crust ‘footprint” undergoing plastic motion, and $\theta_{\rm fp}$ is the co-latitude of the footprint. By asserting that $L_{\rm diss,NS}>25\sigma_{\rm srch}d^{2}{\rm BW}$ for an LPRT at distance $d$ within this model, and using the characteristic $B_{S}\gtrsim(3.2\times 10^{19}\,G)\sqrt{P\dot{P}/2}$, we can solve for the maximum distance $d_{\rm max,NS}(P,\dot{P})$ to which the DN-GPS could detect the LPRT. Figure 10 (top) shows this result with $v_{\rm pl}=10^{4}$ cm yr-1; we see that for $\dot{P}>10^{-13}$ and $P<R^{-1}_{\rm 95}=4.3$ min the search is complete to only 1 kpc. To the edge of the galaxy, approximately 3030\,kpc, the search is only sensitive to $\dot{P}>10^{-10}$. Based on the $\dot{P}$ range of detected LPRTs with $P<R^{-1}_{\rm 95}=4.3$ min, this indicates that the search is not sensitive to magnetar-LPRTs beyond 11\,kpc. In addition, Cooper and Wadiasingh (2024) indicate that the plastic flow velocity can reasonably range from $1<v_{\rm pl}<10^{6}$ cm yr-1, which is proportional to $L_{\rm diss,NS}$ and can significantly alter the horizon distance. Therefore, the search is not complete for magnetar-LPRT objects; we conclude that a population of ultra-long period magnetars may still be present within the Galactic Plane despite our non-detection.
In the Qu and Zhang (2025) White Dwarf-M Dwarf binary model, the companion star orbits within the White Dwarf’s magnetosphere, creating a voltage drop between the two stars via the unipolar inductor effect (Faraday disk). As charge flows between the stars due to the potential drop they create an electrical current, and the magnetosphere dissipates power via Ohmic heating, driving pair creation. Pair charges then radiate via the relativistic electron cyclotron maser mechanism. We adjust the derivation from Qu and Zhang (2025) to allow non-zero spin-down $\dot{P}$; a brief derivation is provided in Appendix A, obtaining the following dissipation luminosity:
@@
Figure 10: $P-\dot{P}$ Constraints from the DN-GPS Survey Using a Neutron Star Model (top; Cooper and Wadiasingh, 2024) and a White Dwarf Model (bottom; Qu and Zhang, 2025). In each plot the thick red lines bound the allowed emission ranges for each model. Vertical grey lines indicate hard limits from the DN-GPS search due to the sampling time and Poissonian burst rate. Pulsars (diamonds), WD binary LPRTs (squares) and all other LPRTs (circles) are colored with their distance estimates, while those without distances are in red. The colorbar also highlights the maximum detectable distance ($d_{\rm max}(P,\dot{P})$) for each model (see Section IV) given the DN-GPS $25\sigma_{\rm srch,P<5\,min}\approx 1030$ mJy and $25\sigma_{\rm srch,P>5\,min}\approx 1200$ mJy detection limits for P<2P<2\,min and $P>2$ min, respectively. Red dashed lines are contours of constant $d_{\rm max}(P,\dot{P})$).
@@ VII Data Availability
The authors would like to thank Adam Dong, Maura McLaughlin, Duncan Lorimer, Scott Ransom, Bing Zhang, Tony Rodriguez, Assaf Horesh, Jayanta Roy, Bhaswati Bhattacharyya, Emmanuel Fonseca, Graham Doskotch, Natasha Hurley-Walker, and Kshitij Bane for creative discussions exploring the unique nature of LPRTs and novel ways to search for them. The authors thank the anonymous referee for offering constructive comments that improved the quality of the final manuscript. This work includes data from the Radio Fundamental Catalogue (RFC; www.doi.org/10.25966/dhrk-zh08) This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2139433. The authors thank staff members of the Owens Valley Radio Observatory and the Caltech radio group, whose tireless efforts were instrumental to the success of the DSA-110. The DSA-110 is supported by the National Science Foundation Mid-Scale Innovations Program in Astronomical Sciences (MSIP) under grant AST-1836018.
@@ Appendix A Derivation of Time-dependent Dissipation Luminosity for the White Dwarf Model
In this appendix, we derive equation 4 incorporating time dependence. Starting from Equation 9 in Qu and Zhang (2025), the total voltage when the White Dwarf and M-dwarf are at a separation $a(t)$ is given by:
@@
Note that in general, we expect $\tau\ll\tau_{\rm age}\sim 1$ Gyr; for example, GPM J1839-10 was detected in archival data over the course of 36 years. Using its upper limit $\dot{P}<3.6\times 10^{-13}$ and measured P=31482.4P=31482.4\,s, we find $\dot{P}\tau/P\sim 10^{-8}$ for $\tau=36\,$years (Horváth et al., 2025b; Men et al., 2025). Even for $\tau=1\,$Gyr, we have $\dot{P}\tau/P\sim 10^{-1}$. Therefore we conclude the solution is valid in general for the LPRT case. Note also that $\zeta$ can vary over a wide range ($\zeta\sim 22$ for GPM J1839-10), but we normalize it to unity for this discussion as in Qu and Zhang (2025).